# AnomalyLMM: Bridging Generative Knowledge and Discriminative Retrieval for Text-Based Person Anomaly Search

### v1

@@ 4.1 Implementation Details
Table 1: Comparison of proposed method with competitive methods on the PAB dataset Yang et al. (2024b). Integrating representative text-to-image retrieval models into the proposed method consistently improves performance. Both CMP and X2VLM benefit from notable gains in Recall@1 (R@1) and mAP when combined with our approach.
@@
   APTM Yang et al. (2023)    72.14    82.78
   IRRA Jiang & Ye (2023)    76.39    86.33
   CLIP Radford et al. (2021)    77.60    87.35
   RaSa Bai et al. (2023)    80.79    89.20
   X-VLM Zeng et al. (2022)    81.95    89.86
   CMP Yang et al. (2024b)    69.51    81.06
@@
Table 2: Ablation studies. (a) Ablation study of primary components, including cloze generation, completion, and comparison & re-ranking. Applying both components yields the best performance. (b) Ablation study of weighting strategies for comparison and re-ranking. The hyper-parameter is the base $\beta$ for exponential decay ($\beta^{n}$) and the step size ($d$) for linear decay. Exponential decay with base $0.5$ achieves the best performance. (c) Ablation study of LLM or LMM for different steps of our method. For cloze generation, QWQ-Max achieves the best results from the LLM group. For cloze completion, QVQ-Max achieves the best results from the LMM group. For comparing & re-ranking, QWQ-Max achieves the best results from the LLM group. (d) Ablation study on the candidate number of the initial ranking list. A length of three achieves the best performance. (e, f) Ablation study of hyper-parameters ($\alpha_{1}$, $\alpha_{2}$) for score balancing. (e) As $\alpha_{1}$ increases from 0.7 to 0.995, while fixing $\alpha_{2}$ as 1, the retrieval performance steadily improves and then saturates at higher values. (f) As $\alpha_{2}$ increases from 0.01 to 0.1, while fixing $\alpha_{1}$ as 1, the performance remains relatively stable, and it drops when $\alpha_{2}$ further increases beyond 0.1.
Method Cloze Generation & Comparison & R@1 mAP Completion Re-Ranking Baseline 83.77 90.43 $w$ Cloze Completion 84.12 90.60 $w$ Re-Ranking 84.22 90.65 Ours 84.73 90.89
@@ 5 Conclusion
Appendix A Appendix
You may include other additional sections here.

### v2

@@ 4.1 Implementation Details
Table 1: Comparison of proposed method with competitive methods on the PAB dataset. Integrating representative text-to-image retrieval models into the proposed method consistently improves performance. Both CMP and X2VLM benefit from notable gains in Recall@1 (R@1) and mAP when combined with our approach.
@@
   APTM    72.14    82.78
   IRRA    76.39    86.33
   CLIP    77.60    87.35
   RaSa    80.79    89.20
   X-VLM    81.95    89.86
   CMP    69.51    81.06
@@
Table 2: Ablation studies. (a) Ablation study of primary components. Applying both yields the best performance. (b) Ablation study of weighting strategies for re-ranking. Exponential decay with base $0.5$ achieves the best performance. (c) Ablation study of LLM or LMM for different steps. (d) Ablation study on the number of initial candidates. A length of three achieves the best performance. (e, f) Ablation study of hyper-parameters ($\alpha_{1}$, $\alpha_{2}$) for score balancing.
Method Cloze Gen. & Comparison & R@1 mAP Completion Re-Ranking Baseline 83.77 90.43 $w$ Cloze Completion 84.12 90.60 $w$ Re-Ranking 84.22 90.65 Ours 84.73 90.89